# TSCMamba: Mamba Meets Multi-View Learning for Time Series Classification

### v1

@@ 4.4 Computational Complexity
In this study, we compared the floating-point operations (FLOPs) of the top-performing methods presented in Table 1. To calculate FLOPs, we set a batch size of 16 across all baselines. For our method, we employed the best-performing hyperparameters, whereas for other baselines, we utilized the recommended parameters specified in the official TimesNet code (Wu et al., 2022a) and Flowformer (Wu et al., 2022b). We leveraged the source code from Ye (2023) to calculate FLOPs. The overall FLOPs, including both forward and backward passes, are presented in Table 2. Notably, our method achieves substantial improvements in terms of FLOPs across all datasets, with the exception of PEMS-SF (PS). This anomaly can be attributed to the projected space ($X$) used to achieve the best result for this dataset, which was set to 1024, thereby impacting the total FLOPs for this dataset only.
@@ 5.1 Component-wise Ablation
Mamba Avg.Pool ROCKET AF EC FD HW HB JV PS SCP1 SCP2 SA UG
62.0 57.0 53.3 74.1 93.0 90.2 92.5 66.7 94.1 93.8
33.1 63.2 34.1 73.2 85.4 81.5 86.7 57.2 74.0 89.1
31.6 64.2 52.0 74.1 94.1 63.0 86.7 60.6 96.7 92.8
31.6 69.4 24.8 76.6 97.0 87.3 91.8 58.3 97.6 86.2
30.0 51.5 49.3 72.7 91.4 84.4 88.7 58.9 90.0 90.3
In this section, we conduct an ablation study to investigate the contribution of individual components in our proposed method. The results are presented in Table 3. A notable observation is that the performance degrades in a large margin across all datasets when the Mamba modules are not utilized, highlighting the importance of incorporating Mamba in our approach. Specifically, when Mamba is not employed (2nd row), the intermediate values are bypassed by scanning operations and directly fed into the DP and MLPs for class logit prediction. In the 3rd row, we present depth-wise max-pooling instead of average-pooling, resulting in an input shape of B,3X. Furthermore, when ROCKET-extracted features are not utilized (4th row), we resort to MLP-extracted features, where the former are non-learnable and the latter are learnable. Additionally, in the 5th row, we explore the effect of replacing additive fusion (AF) with multiplicative fusion (MF), as detailed in Sec.3.3. Notably, while Table 3 largely mirrors the best performance reported in Table  1 across most datasets, the SpokenArabicDigits (SA) dataset exhibits optimal performance when employing depth-wise max-pooling and MLP-based features.
5.2 Hyperparameter Sensitivity
@@
In this section, we discuss the key settings for the Mamba model, as detailed in Gu & Dao (2023). The model operates with four main settings: model dimension (d_model), SSM state expansion factor (d_state), local convolution width (d_conv), and block expansion factor (expand). In our experiments, we automatically set the model dimension based on the input data, while we adjust the other three settings. The importance of fine-tuning these parameters is evident from our tests, shown in Figure 2, which clearly demonstrate their impact on our model’s performance. In addition to Mamba’s hyperparameters, we also tuned the dimension of the projected space ($X$) mentioned in 1.
@@
5.3 Effectiveness of Tango Scanning
Although, our way of scanning (Tango Scanning) at first glance may seem counterintuitive to only forward-based scanning and with another additional reverse flip based scanning (Schiff et al., 2024), it provides substantial improvements in accuracy. This approach, as demonstrated its effectiveness by Figure 3, outperforms traditional forward-scanning and other reverse-flip-based scanning for TSC tasks, making it a valuable strategy for complex scanning scenarios. Our tango scanning is explained in more detail in section 1.
@@ 6 Conclusion and Future Work
We present TSCMamba, an innovative approach for Time Series Classification (TSC) designed to enhance performance while maintaining lower Floating Point Operations (FLOPs). TSCMamba leverages multi-view learning to analyze different views of time-series data, including local and/or global features extracted from time and frequency domains, thereby capturing the essential, discriminative patterns of real-world time-series data. Moreover, the proposed tango scanning mechanism demonstrates TSCMamba’s superiority over conventional scanning methods through extensive experimental validation. Our comprehensive experiments highlight TSCMamba’s exceptional performance in terms of both accuracy and computational efficiency, consistently outperforming current state-of-the-art methods. These results suggest that TSCMamba can serve as a robust and efficient solution for a wide range of TSC applications. Looking ahead, our future work will focus on further enhancing TSCMamba by incorporating self-supervised learning techniques and extending its capabilities to multiple-task learning, in addition to the classification task. We also plan to explore the adaptability of TSCMamba across more diverse and complex time-series datasets, aiming to establish its versatility and robustness in various real-world scenarios. The promising results suggest TSCMamba’s high potential for time series applications.
@@ 7 Acknowledgment
This research is supported in part by the NSF under Grant IIS 2327113 and the NIH under Grants R21AG070909, P30AG072946, and R01HD101508-01. We thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for their support and use of the Lipscomb Compute Cluster and associated research computing resources.

### v2

@@ 4.4 Computational Complexity
In this study, we compared the floating-point operations (FLOPs) (in Table 3) of the top-performing methods (see Table 1). To calculate FLOPs, we set a batch size of 16 across all baselines. For our method, we employed the best-performing hyperparameters, whereas for other baselines, we utilized the recommended parameters specified in the official TimesNet code [45] and Flowformer [43]. We leveraged the source code from [61] to calculate FLOPs. The overall FLOPs, including both forward and backward passes, are presented in Table 3. Notably, our method achieves substantial improvements in terms of FLOPs across all datasets, with the exception of PEMS-SF (PS). This anomaly can be attributed to the projected space ($X$) used to achieve the best result for this dataset, which was set to 1024, thereby impacting the total FLOPs for this dataset only.
@@ 5.1 Component-wise Ablation
We conducted an ablation study to investigate the contribution of individual components in our proposed method. The results are presented in Table 4. Each component is systematically removed or replaced to analyze its effect on classification accuracy.
Performance decreases significantly across all datasets when Mamba modules are removed, suggesting that these modules play a crucial role in our approach’s effectiveness. Specifically, Mamba’s state-space formulation allows it to capture long-range dependencies efficiently, and its removal leads to a notable drop in performance highlighting its importance. When Mamba is not employed (2nd row), the intermediate values are bypassed by scanning operations and directly fed into the DP and MLPs for class logit prediction. This results in significantly reduced model capacity, particularly for datasets requiring long-range contextual understanding.
In the 3rd row, we present depth-wise max-pooling instead of average-pooling, resulting in an input shape of B,3X. The decrease in accuracy across several datasets suggests that average pooling provides better feature aggregation by reducing variance and improving stability in feature representations.
Furthermore, when ROCKET-extracted features are not utilized (4th row), we resort to MLP-extracted features, where the former are non-learnable and the latter are learnable. The substantial accuracy drop indicates that ROCKET’s non-learnable convolutional features provide rich and robust representations that enhance classification. Additionally, in the 5th row, we explore the effect of replacing additive fusion (AF) with multiplicative fusion (MF) (of ROCKET features and spectral features), as detailed in Sec.3.3. While multiplicative fusion offers interaction-based feature integration, the observed performance degradation suggests that additive fusion better preserves the scale and discriminative properties of individual features, leading to more effective multi-view representation learning.
Notably, while Table 4 largely mirrors the best performance reported in Table  1 across most datasets, the SpokenArabicDigits (SA) dataset exhibits optimal performance when employing depth-wise max-pooling and MLP-based features. This suggests that for certain datasets, local feature representations may be more informative than global ones, reinforcing the need for dataset-dependent finetuning.
@@
Mamba Avg.Pool ROCKET AF EC FD HW HB JV PS SCP1 SCP2 SA UG Avg.
62.0 57.0 53.3 74.1 93.0 90.2 92.5 66.7 94.1 93.8 77.67
33.1 63.2 34.1 73.2 85.4 81.5 86.7 57.2 74.0 89.1 67.75
31.6 64.2 52.0 74.1 94.1 63.0 86.7 60.6 96.7 92.8 71.58
31.6 69.4 24.8 76.6 97.0 87.3 91.8 58.3 97.6 86.2 72.06
30.0 51.5 49.3 72.7 91.4 84.4 88.7 58.9 90.0 90.3 70.72
5.2 Comparison with Mamba
In addition to the robust baselines used in Tables 1 and 2, we also evaluated the performance of directly using Mamba [19]. This experiment aims to assess whether our modifications—such as the fusion of spectral and temporal features and the introduction of tango scanning—provide tangible benefits over a standard Mamba implementation.
For the 10 benchmark datasets in Table 1, we implemented our approach that directly applied regular Mamba by processing the MTS data through Mamba modules with a standard scanning protocol. The processed features were then passed through two-stage MLPs for classification, following the architecture in Figure 1. This setup ensures a direct comparison between our TSCMamba approach and the original Mamba framework, isolating the impact of our enhancements.
As shown in Figure 6, TSCMamba achieves superior performance compared to all benchmarks across the 10 datasets, demonstrating that our modifications—particularly tango scanning and multi-view feature fusion—enhance the model’s ability to capture both local and global dependencies in MTS classification. All experiments maintain consistent hyperparameters with those reported in Table 1, ensuring a fair evaluation where improvements can be attributed solely to the architectural enhancements rather than differences in hyperparameter tuning.
Figure 6: Comparison between TSCMamba and standard Mamba when applied to time series data.
5.3 Comparison with BiMamba
In this section, we compare TSCMamba’s channel and token (or time) tango scanning (Eqs. 12 and 11) against BiMamba [27]. To ensure a fair comparison, we maintained the same hyperparameters used to achieve the results in Table 1. Both models were trained under identical settings, including learning rates, batch sizes, and sequence lengths, ensuring that any performance differences arise from the architectural modifications rather than hyperparameter tuning. For BiMamba, we used the official code provided by Schiff et al. [27], with tied weights for the in and out projections.
As illustrated in Figure 7, TSCMamba consistently outperforms BiMamba across all benchmark datasets, demonstrating the effectiveness of tango scanning. The observed improvements indicate that tango scanning more effectively models inter-token dependencies, maximizing pairwise token interactions (For details, please see Subsection 3.6) while maintaining computational efficiency.
Figure 7: Performance comparison in accuracy of TSCMamba with tango scanning versus BiMamba [27] across the benchmark datasets.
5.4 Effectiveness of Tango Scanning
While our tango scanning approach may seem counterintuitive compared to purely forward-based scanning or bidirectional scanning with reverse flips [27], it achieves substantial improvements in accuracy. Unlike standard bidirectional approaches that treat both directions symmetrically, tango scanning strategically employs inversion invariance via flipping the sequence only once, enhancing inter-token interactions while maintaining computational efficiency. This strategy provides more complementary information for time series classification task. We provide more conceptual explanation in Subsection 3.6.
As shown in Figure 8, this approach outperforms both traditional forward scanning and reverse-flip-based scanning methods on TSC tasks, particularly in complex scenarios. This improvement is especially notable in datasets where long-range dependencies are critical, as tango scanning enables more effective information flow across distant tokens.
We provide a theoretical explanation of tango scanning in Section 1. These results highlight the advantages of our approach in effectively modeling sequence-level relationships while avoiding the redundancy and computational overhead typically associated with full bidirectional processing.
6 Hyperparameter Sensitivity
@@
In this section, we discuss the key settings for the Mamba model [19]. The model operates with four main settings: model dimension (d_model), SSM state expansion factor (d_state), local convolution width (d_conv), and block expansion factor (expand). While we derived the model dimension automatically from the input data dimensions, we tuned the remaining three parameters. Figure 9 shows how these parameters affect model performance. In addition to Mamba’s hyperparameters, we also tuned the dimension of the projected space ($X$) mentioned in 1.
@@ 7 Conclusion and Future Work
We present TSCMamba, an innovative approach for Time Series Classification (TSC) that enhances performance while reducing Floating Point Operations (FLOPs). By leveraging multi-view learning, TSCMamba analyzes time-series data through both local and global features extracted from time and frequency domains, effectively and efficiently capturing long-range dependent discriminative patterns in real-world data. Our novel tango scanning mechanism, validated through extensive experiments, demonstrates superior performance compared to conventional scanning methods.
Comprehensive evaluations show that TSCMamba consistently outperforms current state-of-the-art methods in both accuracy and computational efficiency across diverse TSC applications. However, our approach has some limitations. The fusion stage produces a tensor of $BD3X$ dimensions, where the $3X$ factor may increase computational requirements. Additionally, while we currently use CWT for spectral domain features, other transformations might provide complementary information that could further improve performance.
Future work will explore three key directions: incorporating self-supervised learning techniques, extending the framework to multiple-task learning beyond classification, and evaluating TSCMamba’s effectiveness on more complex time-series datasets from various domains.
@@ 8 Acknowledgment
This research is supported in part by the NSF under Grant IIS 2327113 and ITE 2433190 and the NIH under Grants R21AG070909 and P30AG072946. We would like to thank NSF for support for the AI research resource with NAIRR240219. We thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for their support and use of the Lipscomb Compute Cluster and associated research computing resources. We also acknowledge and thank those who created, cleaned, and curated the datasets used in this study.